# Quarter- and half-filled quantum Hall states and their competing interactions in bilayer graphene

### v2

@@ Abstract
Bilayer graphene has emerged as a key platform for non-Abelian fractional quantum Hall (FQH) states, exhibiting multiple half-filled plateaus with large energy gaps. Here, we report four unexpected quarter-filled states and complete the sequence of half-filled plateaus by observing the previously missing $\nu=-\frac{3}{2}$ and $\nu=\frac{1}{2}$ states. Identifying the half-filled plateaus according to their Levin–Halperin daughter states, we reveal an alternating pattern of non-Abelian topological orders. Whenever a pair of $N=1$ Landau levels cross, anti-Pfaffian and Pfaffian develop in the lower and higher levels, respectively. Surprisingly, quarter states occur in $N=0$ levels and are also accompanied by daughters. The mutual exclusion of half- and quarter-filled states indicates a robust competition between the interactions favoring paired states of either two-flux or four-flux composite fermions. Finally, we observe several FQH states that require strong interactions between composite fermions. The systematic pattern of non-Abelian states across two generations of even denominators strengthens the identification of their topological orders and suggests a universal origin.
@@
Composite fermions (CFs), electrons bound to a pair of flux quanta and carrying a charge of $q_{\text{CF}}=(1-2\nu)e$, are pivotal for explaining half-filled states.[24] At half filling, CFs are charge neutral and impervious to magnetic fields. Consequently, they can form a metallic Fermi surface,[10] which is commonly observed in the half-filled zeroth Landau level (0LL). At the same filling, incompressible states develop when the metallic state transitions into a CF superconductor. The attraction necessary for CF pairing arises in the first excited Landau level (1LL), whose single-electron in-plane wave functions differ from those of the 0LL.[25] Alternatively, pairing can be promoted in wide quantum wells where wave functions extend out-of-plane.[26, 27] However, the choice of the pairing channel—and thus, whether the state exhibits Abelian or non-Abelian characteristics—is influenced by subtle effects such as Landau-level mixing. Even more elusive are states formed at quarter fillings, previously reported in GaAs[28, 29, 30, 31, 32] and monolayer graphene.[33] These states comprise charge-neutral pairs of four-flux composite fermions, 4CFs, which condense into a superfluid, similar to the case at half filling.[34, 35]
@@
Attractive interactions at half filling are detrimental to FQH states of the Jain sequence $\nu=\frac{n}{2n+1}$. At these fillings, the CF charge $q_{\text{CF}}$ is non-zero. Consequently, these CFs are sensitive to magnetic fields, and the CF metal exhibits Shubnikov–de Haas (SdH) oscillations, which manifest as incompressible Jain states.[36] When a pairing gap forms on the composite Fermi surface, the SdH oscillations are suppressed, and even lower-order Jain states vanish around half-filled plateaus.[2] A compromise between pairing and adjusting to the magnetic field takes the form of ‘daughter’ states, first proposed by Levin and Halperin.[37] In these states, paired CFs enter bosonic quantum Hall states instead of condensing, resulting in FQH states at distinct filling factors determined by the pairing channel.[38, 39] The unique parent–daughter correspondence enables the identification of topological orders at half filling through straightforward bulk transport measurements.
@@
Bilayer graphene (BLG) has established itself as a versatile platform for studying the superconductivity of both electrons[40] and CFs in even-denominator states.[17, 18, 19, 22] An important feature is the possibility of tuning the Landau level spectrum, spanned by spin, valley, and orbital quantum numbers[41] (see Fig. 1A) and split by interactions,[42, 18, 43] through independently controlling filling factor and displacement field. By exploring this phase space, plateaus at multiple distinct half-integer fillings have been identified according to their daughter states[22, 23, 44]. However, the periodic pattern of non-Abelian states remained hidden due to missing plateaus $\nu=-\frac{3}{2}$ and $\nu=\frac{1}{2}$. Quantum Hall states at quarter-integer fillings and other ‘next-generation’[45, 46, 32] even denominators had not been seen in BLG.
@@
In this work, we report the observation of incompressible FQH states at quarter- and half-filled Landau levels in BLG. First, our measurement of additional plateaus at $\nu=-\frac{3}{2}$ and $\nu=\frac{1}{2}$ reaffirms an orbital index of $N=1$ as a prerequisite to forming incompressible states at half fillings. By observing their daughters, we identify both states as anti-Pfaffian. These plateaus complete the sequence of previously known states and reveal a systematic alternation between anti-Pfaffian and Pfaffian in odd and even Landau levels, respectively. Second, we observe the emergence of incompressible states at four distinct quarter-integer fillings. Surprisingly, all these plateaus arise in $N=0$ levels, which are not expected to favor CF pairing.[47] The topological order of three among the newly discovered quarter states is constrained to the non-Abelian anti-Pfaffian or Abelian anti-(331) by clearly resolved daughter states. The formation of quarter-filled states only at $N=0$, in contrast to half-filled states only at $N=1$, indicates a robust competition between the interactions favoring either paired states of 4CFs or of 2CFs, respectively. Finally, we observed odd and even ‘next-generation’ FQH states at partial fillings of 2CFs.
@@
Results. Our devices (device 1 and device 2) consist of BLG encapsulated by hexagonal boron nitride and graphite layers acting as gates;[48] see Materials and Methods section. Subsequently, we fabricated the heterostructures into a Hall bar, as shown in Supplementary Material (SM) Sec. 1. The electrical transport properties of device 1 were measured in a dilution refrigerator with a base temperature $T=16~{}\mathrm{mK}$ and a magnetic field $B$ up to $18~{}\mathrm{T}$. Device 2 was measured in another dilution refrigerator with a base temperature $T=10~{}\mathrm{mK}$ and a magnetic field up to $12~{}\mathrm{T}$. The data shown in the main manuscript are from device 1 and are consistent with those of device 2 (see SM Sec. 12). Using a standard low-frequency lock-in technique, we measured the longitudinal ($R_{xx}$) and transverse ($R_{xy}$) resistances; see SM Sec. 1. The dual-gate geometry permits independent tuning of filling factor $\nu$ and displacement field $D$, thus granting control over both the orbital ($N=0,1$) and the valley isospin ($K$, $K^{\prime}$).[42, 18]
Fig. 1B shows $R_{xx}$ and $R_{xy}$ as a function of $\nu$ at $D=-80~{}\mathrm{mV/nm}$ and $B=18~{}\mathrm{T}$, where the integer LL has an $N=0$ orbital. We observe a compressible 2CF metal at $\nu=\frac{1}{2}$, flanked by an abundance of SdH oscillations, testifying to the low disorder of the device. As we increase the displacement field to $D=-160~{}\mathrm{mV/nm}$, the Landau level transitions to $N=1$. Here, we observe an incompressible state at $\nu=\frac{1}{2}$ and a strong suppression of the Jain states; see Fig. 1C. The diminished $R_{xx}$ along with a quantized plateau in $R_{xy}$ indicates pairing of 2CFs at $\nu=\frac{1}{2}$. Surprisingly, near $\nu=\frac{3}{4}$, the roles of $N=0$ and $N=1$ are reversed: For $N=0$, we observe a strong suppression of $R_{xx}$ along with a plateau in $R_{xy}$ at $\nu=\frac{3}{4}$, indicating the pairing of 4CFs. By contrast, at $N=1$, this filling hosts a compressible 4CF metal flanked by SdH oscillations. Fig. 1D shows an extended measurement over a continuous range of $D$ which confirms the robustness of both half- and quarter-filled states. Moreover, it shows that these states form irrespective of the valley isospin. Notice that the position of the FQH states changes slightly as either $D$ or $B$ are varied, consistent with previous works.[18, 22] We attribute this effect to a change in quantum capacitance. Subsequently, FQH states were labelled according to their $R_{xy}$ values and the position of their $R_{xx}$ minima with respect to other states. The quarter- and half-filled states at $\nu=\frac{3}{4}$ and $\nu=\frac{1}{2}$ were also observed in device 2, see SM Sec 12.
@@
Figure 1: Interplay between quarter- and half-filled FQH states in BLG. (A) BLG unit cell’s real space single-electron wave functions for the 0LL, eightfold degenerate in valley isospin, orbital, and spin. (B,C) Longitudinal $R_{xx}$ (red) and transverse $R_{xy}$ (blue) resistances as a function of $\nu$ at $T=16~{}\mathrm{mK}$, $B=18~{}\mathrm{T}$ for $D=-80~{}\mathrm{mV/nm}$ ($N=0$) (B) and $D=-160~{}\mathrm{mV/nm}$ ($N=1$) (C). Panels in the top right corners indicate the valley isospin and orbital. The shaded purple (center) and green (sides) regions correspond to FQH states composed of 2CFs and 4CFs. The insets show zoom-ins to $R_{xy}$ plateaus at $\nu=\frac{1}{2}$ (C) and $\nu=\frac{3}{4}$ (B). We attribute these plateaus to pairing of 2CFs and 4CFs, respectively. In (B), $R_{xx}$ data below $\nu=\frac{1}{3}$ were divided by $5$ to accommodate large $R_{xx}$ values. (D) $R_{xx}$ as a function of $\nu$ and $D$, controlled electrostatically by altering the back and top gate voltages, $V_{\mathrm{bg}}$ and $V_{\mathrm{tg}}$. The dotted horizontal lines at $D=-80~{}\mathrm{mV/nm}$ (light blue) and $D=-160~{}\mathrm{mV/nm}$ (orange) indicate the line cuts displayed in (B) and (C), respectively.
We observed all previously reported half-filled states at $\nu=-\frac{5}{2},-\frac{1}{2},\frac{3}{2},\frac{5}{2},\frac{7}{2}$ along with the same daughter states[22]. The new half-filled state at $\nu=-\frac{3}{2}$ emerges in a single pocket around $D=0$, accompanied by daughter states $\nu=-\frac{20}{13},-\frac{25}{17}$. The one at $\nu=\frac{1}{2}$ occurs in two pockets with opposite $D$ polarity, with daughters at $\nu=\frac{6}{13},\frac{9}{17}$. All half-filled states develop only at $N=1$, where the Jain states are strongly suppressed, and quarter-filled states are absent.
To look for additional quarter-filled states, we broaden our study to all eight levels comprising the 0LL. In Fig. 2A, we show $R_{xx}$ as a function of $\nu$ and $D$ for the hole-doped (left) and electron-doped (right) sides. Dark-blue areas mark $R_{xx}<10~{}\Omega$ where the bulk becomes incompressible. On the hole (electron) side, the $R_{xx}$ measurement is overshadowed by artifacts for negative (positive) $D$ due to the contact’s fabrication scheme; see Materials and Methods section.
Figure 2: Even-denominator FQH states in the 0LL of BLG. (A) $R_{xx}$ as a function of $\nu$ and $D$ for $-4\leq\nu<0$ (left panel) and $0<\nu\leq 4$ (right panel). Areas highlighted by dashed white lines refer to high-resolution measurements in (C-F) and Fig. 5. See SM Sec. 4 for enlarged view. (B) Schematic summary of all even- and selected odd-denominator FQH states observed in (A), with higher-order Jain states omitted to increase visibility. Vertical lines mark observed FQH states; bold vertical lines mark newly observed FQH states. Dashed lines mark valley and orbital crossings along the ($\nu$, $D$) phase space. Green and purple labels indicate Pfaffian (Pf) and anti-Pfaffian (aPf) states, respectively, according to their Levin–Halperin daughters. (C-F) $R_{xx}$ measurements near the four quarter-filled states. White dots in (E) mark the values of $D$ where we measured the gap $\Delta_{\frac{3}{4}}$ in Fig 4E.
Upon tuning the orbital index from $N=1$ to $N=0$, we observed quarter-filled states at $\nu=\frac{3}{4}+(-4,-2,0,+2)$ and a suppression of the 4CF Jain sequences around these fillings. The systematic occurrence of these states at every second level suggests a common origin. Additionally, we find daughter states at $\nu=\frac{9}{25}+(-4,-2,0)$ but not at $\nu=\frac{9}{25}+2$; see Fig. 2C–F.
@@
In Fig. 2B, we summarize the observed quarter- and half-filled states along with their daughter states, highlighting their dependence on the orbital and valley isospin indices. The newly found half- and quarter-filled states are shown by bold, solid black lines. We already labeled the even-denominator states according to their identification via Levin–Halperin daughter states and in-plane $B$ measurements, which we discuss below.
@@
We proceed with a detailed characterization of all half-filled states. To avoid ambiguity, we identify all states according to the topological phase of the electrons for either sign of $\nu$. In Fig. 3A, we show the filling factors of daughters associated with different pairing channels of 2CFs.[38, 39] In Fig. 3B, we show $R_{xx}$ and $R_{xy}$ around $\nu=\frac{1}{2}$ at $D=-140~{}\mathrm{mV/nm}$, with strong dips in $R_{xx}$ flanking the plateau at half filling. The absence of higher-order Jain states beyond $\nu=\frac{2}{5},\frac{3}{5}$ suggests that the dips near $\nu=\frac{1}{2}$ are daughter states. We mark possible daughter-state fillings by vertical lines for the most relevant pairing channels, including Moore–Read and anti-Pfaffian favored by numerics[3] and PH-Pfaffian indicated by thermal measurements in GaAs.[49, 50, 51] The observed dips coincide with $\frac{6}{13}$ and $\frac{9}{17}$, supporting anti-Pfaffian pairing for $\nu=\frac{1}{2}$. Our analysis of the other half-filled states and the plateaus of their daughters, which agrees with previous reports,[22] is shown in SM Sec. 8–10. The identification of paired states based on their daughters is indirect and has not been independently confirmed by direct measurements, such as thermal conductance[49, 50] or upstream noise at interfaces[51]. Still, the periodic pattern of topological orders indicated by daughters[38, 39] and its systematic association with crossing $N=1$ levels provides strong support for its veracity.
@@
Fig. 3C shows the dependence of the gap $\Delta_{\nu=\frac{1}{2}}$ on an in-plane magnetic field $B_{\parallel}$ at an out-of-plane field of $B_{\perp}=15~{}\mathrm{T}$ (see SM Sec. 7 for more details). We observe only negligible variations of $\Delta_{\frac{1}{2}}$ up to the largest accessible value of $B_{\parallel}=10~{}\mathrm{T}$ and conclude that the $\nu=\frac{1}{2}$ state is fully spin-polarized, as expected for the anti-Pfaffian. A detailed characterization of $\nu=-\frac{3}{2}$ also indicates a spin-polarized state; see SM Sec. 10.
Figure 3: Topological orders of half-filled states in BLG. (A) Classification of half-filled states and their identification based on daughter-state fillings. The pairing channel $\ell$ corresponds to the number of Majorana edge channels (see SM Sec. 13). (B) $R_{xx}$ and $R_{xy}$ around $\nu=\frac{1}{2}$, measured at $B=18~{}\mathrm{T}$ and $D=-140~{}\mathrm{mV/nm}$, along with possible daughter states marked by vertical dashed lines. (C) The thermal activation gap $\Delta_{\frac{1}{2}}$ measured as a function of $B_{\parallel}$ at $B_{\perp}=15~{}\mathrm{T}$. (D) $R_{xx}$ around $\nu=\frac{3}{2}$ measured as a function of $B$ for $D=-140~{}\mathrm{mV/nm}$. (E) Identification of all observed half-filled states based on their daughters, along with the measured thermal activation gaps. The gap of all the half-filled states was measured at $B=18~{}\mathrm{T}$, except for $\nu=-\frac{3}{2}$ which was measured at $B=11.2~{}\mathrm{T}$.
@@
We summarize our identification of the half-filled states in the table of Fig. 3E (see SM Secs. 5 and 8 for gap measurements and daughters of other half-filled states). Strikingly, all observed anti-Pfaffians (aPf) occur in a pair with a Pfaffian (Pf) in the vicinity of crossing $N=1$ levels. There is one such crossing for ($-\frac{3}{2}$,$-\frac{1}{2}$), two for ($\frac{1}{2}$,$\frac{3}{2}$), and one for ($\frac{5}{2}$,$\frac{7}{2}$), each exhibiting (aPf, Pf) on both sides of the crossing. This observation suggests that Landau-level mixing within these pairs is most relevant for determining the topological order at half filling. Pfaffian or anti-Pfaffian are favored by Landau-level mixing,[52, 53, 54] which is opposite for the two members of each pair. In the lower one, electrons can virtually occupy the empty level above; in the upper one, holes can virtually occupy the filled level below. By contrast, the spin and valley isospin appear to play subordinate roles. The largest gap occured at $\nu=\frac{3}{2}$. Fig. 3D shows $R_{xx}$ near this filling as a function of $\nu$ and $B$. The incompressible state at $\nu=\frac{3}{2}$ persists down to $B=6~{}\mathrm{T}$, which places half-filled states in BLG within reach of commonly available magnetic-field setups.
We turn to a detailed analysis of the quarter-filled states. Similar to their cousins at half filling, quarter states can also exhibit characteristic daughters.[38] In Fig. 4A, we show the filling factors of daughters associated with the main pairing channels. In Fig. 4B, we show $R_{xx}$ and $R_{xy}$ around $\nu=\frac{3}{4}$ with the possible daughter states indicated by dashed lines.
The state at $\nu=\frac{19}{25}$ shows remarkable agreement with our data, thereby constraining the $\nu=\frac{3}{4}$ topological order to either the non-Abelian anti-Pfaffian or the Abelian anti-(331). To determine the spin polarization of the $\nu=\frac{3}{4}$ state, we measured its gap while varying $B_{\parallel}$ at $B_{\perp}=15~{}\mathrm{T}$; see Fig. 4C and SM Sec. 7 for more details. The weak dependence of $\Delta_{\frac{3}{4}}$ on $B_{\parallel}$ indicates a fully spin-polarized state. In conjunction with the observed daughter state, this observation supports the non-Abelian anti-Pfaffian ground state, as the spin-singlet anti-(331) state should not be spin-polarized.[55] However, a valley-singlet anti-(331) state remains possible.

### v3

# Quarter- and half-filled quantum Hall states and their topological orders revealed by daughter states in bilayer graphene

@@ Abstract
Even-denominator fractional quantum Hall states are promising candidates for fault-tolerant quantum computing due to their underlying non-Abelian topological order. However, the topological order of these states remains hotly debated. Here, we report transport measurements on ultra-clean bilayer graphene heterostructures, where we observed four quarter-filled states and their corresponding Levin–Halperin daughter states, constraining their topological order. Moreover, we complete the sequence of half-filled plateaus by detecting states at $\nu=-\frac{3}{2}$ and $\nu=\frac{1}{2}$ whose daughters suggest an alternating sequence of non-Abelian orders. This pattern suggests a universal origin supporting their use in identifying topological order at even-denominator fillings, though further confirmation is needed via direct measurements. The observed quarter- and half-filled states appear in $N=0$ and $N=1$ Landau levels, respectively, and thus highlight a competition between interactions favoring paired states of either four- or two-flux composite fermions. Additionally, we observe several ‘next-generation’ quantum Hall states that require strong interactions between composite fermions.
@@
Composite fermions (CFs), electrons bound to a pair of flux quanta and carrying a charge of $q_{\text{CF}}=(1-2\nu)e$, are pivotal for explaining half-filled states.[24] At half filling, CFs are charge neutral and impervious to magnetic fields. Consequently, they can form a metallic Fermi surface,[10] which is commonly observed in the half-filled zeroth Landau level (LL), denoted by $N=0$. At the same filling, incompressible states develop when the metallic state transitions into a CF superconductor. The attraction necessary for CF pairing arises in the first excited LL ($N=1$), whose single-electron in-plane wave functions differ from those at N=0N=0.[25] Alternatively, pairing can be promoted in wide quantum wells where wave functions extend out-of-plane.[26, 27] However, the choice of the pairing channel—and thus, whether the state exhibits Abelian or non-Abelian characteristics—is influenced by subtle effects such as Landau-level mixing (LLM). Even more elusive are states formed at quarter fillings, previously reported via magneto-resistance measurements on GaAs[28, 29, 30, 31, 32] and in a capacitance measurement on monolayer graphene.[33] These states comprise charge-neutral pairs of four-flux composite fermions, 4CFs, which condense into a superfluid, similar to the case at half filling.[34, 35]
@@
Attractive interactions at half filling are detrimental to FQH states of the Jain sequence at $\nu=\frac{n}{2n\pm 1}$.[16] At these fillings, the CF charge $q_{\text{CF}}$ is non-zero. Consequently, these CFs are sensitive to magnetic fields, and the CF metal exhibits Shubnikov–de Haas (SdH) oscillations, which manifest as incompressible Jain states.[36] When a pairing gap forms on the composite Fermi surface, the SdH oscillations are suppressed, and even lower-order Jain states vanish around half-filled plateaus.[2, 16] A compromise between pairing and adjusting to the magnetic field takes the form of ‘daughter’ states, first proposed by Levin and Halperin.[37] In these states, paired CFs enter bosonic quantum Hall states instead of condensing, resulting in FQH states at distinct filling factors determined by the pairing channel.[38, 39] This unique parent–daughter correspondence may shed light on the topological order at half filling and was used to identify states in bilayer graphene (BLG)[22, 40] and GaAs.[41, 16] However, this approach warrants confirmation through direct observations at half filling, such as via thermal conductance measurements.
@@
BLG has established itself as a versatile platform for studying the superconductivity of both electrons[42] and CFs in even-denominator states.[17, 18, 19, 22] In contrast to GaAs, where $N=0$ and $N=1$ are separated by the cyclotron energy, these levels are degenerate in BLG. Here, the single-particle electron wavefunction is confined to a single layer when $N=0$, whereas, for $N=1$ it extends across both layers (Fig. 1A, left), reducing Coulomb repulsion and facilitating electron pairing.[43] This degeneracy, combined with additional degeneracies in spin ($\sigma=\uparrow,\downarrow$) and valley ($K,K^{\prime}$) quantum numbers, gives rise to a total of 8 degenerate ground states. In high-mobility heterostructures measured at low temperatures and high magnetic fields, these degeneracies are lifted, revealing a phase space spanned by the displacement field and filling factor, with 16 crossings of integer LLs. (Fig. 1A, right).[44, 18, 45, 46] By exploring this phase space, plateaus at multiple distinct half-integer fillings have been identified as either Pfaffian or anti-Pfaffian according to their daughter states[22, 40]. However, the periodic pattern of non-Abelian states remained hidden due to missing plateaus at $\nu=-\frac{3}{2}$ and $\nu=\frac{1}{2}$. Quantum Hall states at quarter-integer fillings and other ‘next-generation’[47, 48, 32] even denominators had not been seen in BLG, to the best of our knowledge.
@@
In this work, we report the observation of incompressible FQH states at quarter- and half-filled LLs in BLG, which we attribute to a combination of high sample quality, highly transparent contacts, and an ultra-low noise measurement setup (see Methods section). First, our measurement of additional plateaus at $\nu=-\frac{3}{2}$ and $\nu=\frac{1}{2}$ reaffirms an orbital index of $N=1$ as a prerequisite to forming incompressible states at half fillings. By observing their daughters, we identify both states as anti-Pfaffian. These plateaus complete the sequence of previously known states and reveal a systematic alternation between anti-Pfaffian and Pfaffian in odd and even LLs, respectively. Moreover, our work directly links the half-filled plateaus and their daughters to Landau level crossings, which introduced a perspective that was not addressed in previous works. Second, we observe the emergence of incompressible states at four distinct quarter-integer fillings, with plateaus appearing exclusively within the $N=0$ level, consistent with previous studies.[28, 29, 30, 31, 32]. Moreover, by resolving their daughter states, we were able to constrain their topological order. The formation of quarter-filled states at $N=0$, in contrast to half-filled states at $N=1$, indicates a competition between the interactions favoring either paired states of 4CFs or of 2CFs, respectively. Finally, we observed odd and even ‘next-generation’ FQH states at partial fillings of 2CFs.
@@
Results. Our devices (device 1 and device 2) consist of BLG encapsulated by hexagonal boron nitride and graphite layers acting as gates;[49] see Methods section. Subsequently, we fabricated the heterostructures into a Hall bar, as shown in supplementary note 1. The electrical transport properties of device 1 were measured in a dilution refrigerator with a base temperature $T=16~\mathrm{mK}$ and a magnetic field $B$ up to $18~\mathrm{T}$. Device 2 was measured in another dilution refrigerator with a base temperature $T=10~\mathrm{mK}$ and a magnetic field up to $12~\mathrm{T}$. The data shown in the main manuscript are from device 1 and are consistent with those of device 2 (see supplementary note 12). Using a standard low-frequency lock-in technique, we measured the longitudinal ($R_{xx}$) and transverse ($R_{xy}$) resistances; see supplementary note 1. Subsequently, fully developed FQH states were identified according to their $R_{xy}$ values and accompanying minima in $R_{xx}$. For less developed plateaus, the filling factors were determined by the position of their $R_{xx}$ minima with respect to fully developed plateaus, e.g., $\frac{2}{5}$, $\frac{1}{2}$, and $\frac{3}{5}$, etc. The dual-gate geometry permits independent tuning of filling factor $\nu$ and displacement field $D$, thus granting control over both the orbital ($N=0,1$) and the valley isospin ($K$, $K^{\prime}$).[44, 18]
[…]
@@
Figure 1: Interplay between quarter- and half-filled FQH states in BLG. (A) BLG unit cell’s real space single-electron wave functions for the zero-energy $L$, eightfold degenerate in valley isospin, orbital, and spin (left) and the evolution of $L$s as their degeneracy is lifted (right). Dotted boxes refer to crossings of $N=1$ $L$s. The solid black line box refers to level crossings observed in panel D. (B,C) Longitudinal $R_{xx}$ (red) and transverse $R_{xy}$ (blue) resistances as a function of $\nu$ at $T=16~\mathrm{mK}$, $B=18~\mathrm{T}$ for $D=-80~\mathrm{mV/nm}$ ($N=0$) (B) and $D=-160~\mathrm{mV/nm}$ ($N=1$) (C). Panels in the top right corners indicate the valley isospin and orbital. The shaded purple (center) and green (sides) regions correspond to FQH states composed of 2CFs and 4CFs. The insets show zoom-ins to $R_{xy}$ plateaus at $\nu=\frac{1}{2}$ (C) and $\nu=\frac{3}{4}$ (B). We attribute these plateaus to pairing of 2CFs and 4CFs, respectively. In (B), $R_{xx}$ data below $\nu=\frac{1}{3}$ were divided by 55 to accommodate large $R_{xx}$ values. (D) $R_{xx}$ as a function of $\nu$ and $D$, controlled electrostatically by altering the back and top gate voltages, $V_{\mathrm{bg}}$ and $V_{\mathrm{tg}}$. The dotted horizontal lines at $D=-80~\mathrm{mV/nm}$ (light blue) and $D=-160~\mathrm{mV/nm}$ (orange) indicate the line cuts displayed in (B) and (C), respectively.
We now broaden our study to all eight levels comprising the zero-energy LL. In Fig. 2A, we show $R_{xx}$ as a function of $\nu$ and $D$ for the hole-doped (left) and electron-doped (right) sides. Dark-blue areas mark $R_{xx}<10~\Omega$ where the bulk becomes incompressible. On the hole (electron) side, the $R_{xx}$ measurement is overshadowed by artifacts for negative (positive) $D$ due to the contact fabrication scheme; see Methods section. We observed all previously reported half-filled states at $\nu=-\frac{5}{2},-\frac{1}{2},\frac{3}{2},\frac{5}{2},\frac{7}{2}$ along with the same daughter states[22]. All theoretically predicted daughter states occur at filling factors that coincide with the Jain sequence $\nu_{\text{Jain}}=\frac{n}{2n+1}$ for relatively large $|n|$. However, the gaps of the Jain sequence are expected to decrease as $\Delta\sim\frac{1}{2n+1}$ toward half filling, consistent with observations in $N=0$ LLs.[50, 16] We identify anomalously strong states that violate this pattern as daughter states, by comparing the strength of the $R_{xx}$ minima. For example, in Fig. 1C, we observe Jain states only up to $n=3$ ($\nu=\frac{3}{7}$) on the particle-like side and down to n=4n=-4 ($\nu=\frac{4}{7}$) on the hole-conjugate side. This observation strongly suggests that the prominent states at $n=6$ ($\nu=\frac{6}{13}$) and n=9n=-9 ($\nu=\frac{9}{17}$) have a fundamentally different character, i.e., they are daughter states.
The half-filled state at $\nu=-\frac{3}{2}$ emerges in a single pocket around $D=0$, accompanied by daughter states at $\nu=-\frac{20}{13},-\frac{25}{17}$ (see supplementary note 10 Fig. 21). The one at $\nu=\frac{1}{2}$ occurs in two pockets centered around $D=\pm 170~\mathrm{mV/nm}$, with daughters at $\nu=\frac{6}{13},\frac{9}{17}$. All half-filled states develop only at $N=1$, where the Jain states are strongly suppressed, and quarter-filled states are absent. Upon tuning the orbital index from $N=1$ to $N=0$, we observed quarter-filled states at $\nu=\frac{3}{4}+(-4,-2,0,+2)$ and a suppression of the 4CF Jain sequences around these fillings. This observation is consistent with previous studies in GaAs.[29, 28, 30, 32, 51] Here, quarter-filled states in hole-doped systems were attributed to LLM, estimated to be $\frac{E_{\mathrm{C}}}{\Delta}\approx 3-8$, where $E_{C}$ is the Coulomb energy and $\Delta$ is the cyclotron energy. The estimated LLM in BLG is similar, i.e., $\frac{E_{\mathrm{C}}}{\Delta_{01}}\approx 5-10$, with $\Delta_{01}$ the gap between the $N=0,1$ levels of the same spin and valley,[19] and could therefore promote 4CF pairing. More importantly, we observe their daughter states at $\nu=\frac{9}{25}+(-4,-2,0)$; see Fig. 2C–F, which we use to constrain their topological order. The systematic appearance of these states in every second LL in BLG along with identical daughters suggests a common origin.
@@
In Fig. 2B, we summarize the observed quarter- and half-filled states along with their daughter states, highlighting their dependence on the orbital and valley isospin indices. The half- and quarter-filled states that were not observed in previous works[18, 19, 22], are shown by bold, solid black lines; see supplementary Figs. 11 and 12 for an enlarged view of hole and electron side along with their schematics. We already labeled the even-denominator states according to their identification via Levin–Halperin daughter states and in-plane $B$ measurements, which we discuss below.
[…]
@@
We proceed with a detailed characterization of all half-filled states. To avoid ambiguity, we identify all states according to the topological phase of the electrons for either sign of $\nu$. In Fig. 3A, we show the filling factors of daughters associated with different pairing channels of 2CFs.[38, 39] In Fig. 3B, we show $R_{xx}$ and $R_{xy}$ around $\nu=\frac{1}{2}$ at $D=-140~\mathrm{mV/nm}$, with strong dips in $R_{xx}$ flanking the plateau at half filling. The absence of higher-order Jain states beyond $\nu=\frac{2}{5},\frac{3}{5}$ suggests that the dips near $\nu=\frac{1}{2}$ are daughter states. We mark possible daughter-state fillings by vertical dashed lines with their colors referring to the topological order of the parent state from Fig. 3A, including Pfaffian and anti-Pfaffian favored by numerics[3] and PH-Pfaffian suggested by thermal conductance measurements in GaAs.[52, 53, 54] The observed dips coincide with $\frac{6}{13}$ and $\frac{9}{17}$ (see supplementary note 9 for quantization of daughters), supporting anti-Pfaffian pairing for $\nu=\frac{1}{2}$. Similarly, based on the daughters, we identified the state at $\nu=-\frac{3}{2}$ to be anti-Pfaffian (supplementary note 10). Our analysis of the other half-filled states and the plateaus of their daughters, which agrees with previous reports,[22] is shown in supplementary notes 8 and 9. The identification of paired states based on their daughters is indirect and has not been independently confirmed by direct measurements, such as thermal conductance[52, 53] or upstream noise at interfaces.[54] Still, the periodic pattern of topological orders indicated by daughters[38, 39] and its systematic association with crossing $N=1$ levels (see dotted boxes in Fig. 1A) provides strong support for its veracity.
@@
Fig. 3C shows the dependence of the activation gap $\Delta_{\nu=\frac{1}{2}}$ on an in-plane magnetic field $B_{\parallel}$ at an out-of-plane field of $B_{\perp}=15~\mathrm{T}$. The activation gap of the states examined in the manuscript is determined from the slope of the linear fit to the Arrhenius plot of $R_{xx}$ versus temperature. The error bars reflect the uncertainty in this linear fit (see supplementary note  7 for more details). This method of determining the activation gap and its associated error bar is applied to all figures and tables in the manuscript. We observe negligible variations of $\Delta_{\frac{1}{2}}$ up to the largest accessible value of $B_{\parallel}=10~\mathrm{T}$ which corresponds to a Zeeman energy of approximately $14~\mathrm{K}$, more than an order of magnitude larger than $\Delta_{\frac{1}{2}}$, similar to previous work.[18] We, therefore, conclude that the $\nu=\frac{1}{2}$ state is fully spin-polarized, as expected for the anti-Pfaffian. Note that, to fully validate the spin polarization, a direct approach such as NMR study is required.[55, 56] A detailed characterization of $\nu=-\frac{3}{2}$ also indicates a spin-polarized state; see supplementary note 10.
@@
We summarize our identification of the half-filled states in the table of Fig. 3D (see supplementary notes 5 and 8 for gap measurements and daughters of other half-filled states). Strikingly, all observed anti-Pfaffians (aPf) occur in a pair with a Pfaffian (Pf) in the vicinity of two crossing $N=1$ levels. There is one such crossing for ($\frac{3}{2}$$\frac{1}{2}$), two for ($\frac{1}{2}$$\frac{3}{2}$), and one for ($\frac{5}{2}$$\frac{7}{2}$), each exhibiting (aPf, Pf) on both sides of the crossing (see Fig. 1A and supplementary note 1 Fig. 2 for an illustration). This observation suggests that LLM within these pairs is most relevant for determining the topological order at half filling. Pfaffian or anti-Pfaffian are favored by LLM,[57, 58, 59] which is opposite for the two members of each pair. In the lower one, electrons can virtually occupy the empty level above; in the upper one, holes can virtually occupy the filled level below. By contrast, the spin and valley isospin appear to play subordinate roles. The largest gap occurred at $\nu=\frac{3}{2}$. Fig. 3E shows $R_{xx}$ around this filling as a function of $\nu$ and $B$. The incompressible state at $\nu=\frac{3}{2}$ persists down to $B=6~\mathrm{T}$, which places half-filled states in BLG within reach of commonly available magnetic-field setups.